Hailey15:

math help A normal random variable X has mean 40 and standard deviation 16. Find the area under the normal curve above the interval 16-60. 0.7324 0.8962 0.8276 0.8089

dude:

How many standard deviations is 16 and 60 from the mean?

Hailey15:

no idea

dude:

You have to find the z-score Do you know how to find z-score?

Hailey15:

no I dont :(

dude:

\(z=\large\frac{x-\mu}{\sigma}\) \(\mu\)=mean \(\sigma\) = Standard deviation x = your value (In this case, its 16 and 60)

Hailey15:

oh I see

dude:

So I'll do the first one \(z=\large \frac{16-40}{16}\) \(z=-1.5\) standard deviations from the mean

Hailey15:

so the second is 1.25?

dude:

\(z=\large \frac{60-40}{16}\) \(z=1.25\) Right

Hailey15:

then what do I do?

babygirlz666:

i am so confused

babygirlz666:

this isnt helping

Hailey15:

He is doing a great job :) I can understand I just don't know what to do next lol

dude:

For this you need to use the z-score table

1 attachment
dude:

Sorry that response took so long btw

Hailey15:

its fine :) and if you dont mind can you please show me how to use it

dude:

Okay so the left shows the z-score, the top part is if there are more than 1 numbers after the decimal Ex. -1.5 is 0.668 but -1.52 is -.643 RSKlavQlQbGlQ_e_QBtajA.png

dude:

Correction: -1.52 is 0.643*

Hailey15:

oh so we ignore 0.668?

dude:

No no, we use 0.668 because we had -1.5 as our z-score before, the -1.52 was just an example on how to use it

dude:

This is the positive z-score values 8573955.png?759

dude:

^We're missing 1.25, so do you know how to find it using that table?

Hailey15:

is it .8944?

dude:

Right, so the issue with this is that is includes all of the values from 0-1.25, in order to get rid of this, we subtract 0.8944-0.0668 (What we got from the z-score of -1.5 and the z-score of 1.25)

Hailey15:

right

dude:

And that should be your answer

Hailey15:

ooooh I seeeeee omg I finally get it!! Thank you sooo much! I appreciate it :)

dude:

I felt like I explained this too quickly, if you want me to review another question I could try to re-explain it

Hailey15:

That would be great! omg your thank you :)

dude:

Yeh In a nutshell, ______________________________________________________________________ Find the z-score using the equation \(z=\large \frac{x-\mu}{\sigma}\) Then use the tables I gave you to find the % area (Careful for the positive and negative on both) Subtract the highest z-score value from the lowest z-score and that should give you your answer ______________________________________________________________________

Hailey15:

Ok this is a similar one A normal distribution has mean 200 and standard deviation 50. Find the area under the normal curve from the mean to 224

dude:

Make a new post before this gets laggy :P

Hailey15:

oh ok

Hailey15:

I just wanted to say thank you so much for this! I came back to review for my finals and this really helped!

dude:

Haha no problem